# Learning More Effective Representations for Dense Retrieval through Deliberate Thinking Before Search

### v1

@@ 4 Experimental Methodology
Model Size ($\rightarrow$) / 4.8B 5.8B 8B 7B 7B 7B 2.4B 4B
@@
Table 1: Overall Retrieval Performances on BEIR Benchmark. indicates the 11 most representative BEIR tasks used in CPT Neelakantan et al. (2022) and Avg CPT Sub reflects the average performance across these tasks.
Datasets. We train all Debater models using the public portion of the E5 dataset Wang et al. (2024a); Springer et al. (2025), which comprises approximately 1.5M samples. The retrieval effectiveness of Debater is evaluated on the BEIR benchmark Thakur et al. (2021), which includes 18 datasets that span a variety of domains. Our evaluation focuses on the 14 publicly available datasets used for the zero-shot retrieval task. Additional details about the training dataset and evaluation benchmark can be found in Appendix 6.
Evaluation Metrics. To evaluate the retrieval effectiveness of Debater, we use nDCG@10, the standard metric for the BEIR benchmark. The metric implementation follows the pytrec-eval toolkit Gysel and de Rijke (2018), which is consistent with prior work Zhu et al. (2023).
Baselines. We compare Debater with several baseline retrievers implemented with different language models. GTR Ni et al. (2022) employs large dual encoder-only models to build a dense retriever, while SGPT Muennighoff (2022) trains dense retrieval models using decoder-only architectures. Emb-V3111https://cohere.com/blog/introducing-embed-v3 is a commercial text retrieval model provided by Cohere. PromptReps Zhuang et al. (2024) directly prompts LLMs to generate dense representations without supervision. RepLLaMA Ma et al. (2024) and E5-Mistral Wang et al. (2024b) fine-tune LLMs as dense retrievers, using the hidden state of an additional end-of-sequence token to represent the input context. Notably, E5-Mistral is trained on the same dataset as Debater but leverages a larger foundational model.
Implementation Details. We initialize the Debater models with MiniCPM-2.4B and MiniCPM-4B Hu et al. (2024). All Debater models are trained for 1,000 steps using the AdamW optimizer with a batch size of 256. The learning rate follows a cosine decay schedule, with a warm-up phase covering the first 3% of the total iterations, peaking at 2e-4. We train Debater using hybrid negatives, including one hard negative from the E5 dataset and seven in-batch negatives. The CoD length for all Debater models is set to 8. Debater is implemented using the OpenMatch toolkit Yu et al. (2023), with flash-attention Dao et al. (2022) and LoRA Hu et al. (2022) enabled to mitigate memory constraints and improve computational efficiency.
5 Evaluation Results
@@
[…]
5.1 Overall Performance
The overall performance of Debater and the baseline retrievers is shown in Table 1. Additional experimental results comparing Debater with other baseline retrievers are provided in Appendix A.4.
5.2 Ablation Study
@@
5.3 Effectiveness of Chain-of-Deliberation with Different Thinking Depths
@@
Figure 3: Retrieval Performance of Debater with Different Thinking Depths. We set the length of the CoD to train different Debater models and evaluate them on different subsets of BEIR.
@@
Figure 4: Performance of Debater at Different Thinking Steps. We collect all documents from each thinking step to demonstrate the retrieval performance of Debater across different stages of reasoning.
5.4 Retrieval Performance of CoD-Generated Document Representations
@@
As shown in Figure 4, during the early stages of reasoning (e.g., steps 1 and 2), retrieval performance is relatively low and even decreases. This suggests that initial embeddings, based on minimal deliberation, may lack the nuanced understanding required for effective retrieval. However, as the number of thinking steps increases, performance generally improves, indicating that more deliberation leads to more refined embeddings for retrieval tasks. These improvements demonstrate that Debater helps LLMs better capture relevant information from the tokens of documents and previous decoded representations in the thinking steps. On the other hand, Debater’s performance gradually reaches a plateau as the thinking steps increase, suggesting that after a certain point, the embeddings become sufficiently fine-grained, and further deliberation may result in limited improvements and unnecessary computational complexity.
5.5 Characteristics of the Embeddings Generated by Debater
@@
Figure 5: Similarity Relationship Between Adjacent Position Embeddings. Both Debater w/o SD and Debater are compared in our experiments. Darker blue indicates a higher similarity score.
@@
Figure 6: Similarity Scores Between the First Seven Embeddings and the Last Embedding. The last embedding is used as the representation of documents for retrieval.
@@
Contributions of CoD Steps to Document Representations. Figure 6 illustrates the similarity relationship between embeddings at intermediate thinking steps of CoD and the final document representation generated at the last thinking step. This helps us explore the contributions of different thinking steps to the final document representations.
@@
6 Conclusion
@@
Limitation
Debater demonstrates its effectiveness in activating the reasoning capabilities of LLMs to enhance document representation. However, each thinking embedding requires interaction with the query embedding during the training of Debater, which introduces additional memory costs for GPU. Furthermore, we only employ 2$\times$ NVIDIA A100-40G GPU to train and evaluate our Debater models. Due to the high embedding dimension of the LLM-based retriever, it takes us over a week to build the index for the BEIR evaluation using Debater-2.4B, and over 2 weeks on Debater-4B. Thus, we do not further explore larger Debater models (e.g., Debater-7B) in this paper due to limitations in computational resources. However, our experimental results indicate that the Debater-4B model outperforms all 7B-scale baseline models, highlighting its effectiveness.
Appendix A Appendix
A.1 Licenses
The E5 dataset and BEIR benchmark are released under the Apache License 2.0. The terms of use can be found on their Github pages. This license allows users to freely use, modify, and distribute the data, permitting academic use of their dataset.
A.2 More Details of Datasets Used in Experiments
In this section, we provide a detailed description of the training dataset and the evaluation benchmark for our Debater.
E5 Dataset. We employ the E5 dataset Wang et al. (2024a); Springer et al. (2025) for training, comprising a collection of publicly available datasets. The E5 dataset is carefully curated to encompass a wide range of retrieval scenarios and tasks. We present the statistics of the E5 dataset in Table 3. The instructions used in the E5 dataset are shown in Table 5. The full dataset can be available at their website [^2] https://github.com/jakespringer/echo-embeddings.
BEIR Benchmark. We evaluate our Debater models using the BEIR benchmark, which includes 18 datasets from 9 heterogeneous retrieval tasks. We focus on the 14 publicly available datasets and use the standard instructions in MTEB Muennighoff et al. (2023). The statistics for these datasets are provided in Table 4, and the instructions can be found in Table 6.
Dataset #Samples Proportion
ELI5 Fan et al. (2019) 32,547 2.16%
HotpotQA Yang et al. (2018) 90,447 5.99%
FEVER Thorne et al. (2018) 101,578 6.73%
MIRACL Zhang et al. (2023b) 32,561 2.16%
MSMARCO Passage Ranking Bajaj et al. (2016) 249,592 16.53%
MSMARCO Document Ranking Bajaj et al. (2016) 73,400 4.86%
NQ Kwiatkowski et al. (2019) 100,231 6.64%
NLI Gao et al. (2021) 277,230 18.36%
SQuAD Rajpurkar et al. (2016) 87,599 5.80%
TriviaQA Joshi et al. (2017) 73,346 4.86%
Quora Duplicate Questions DataCanary et al. (2017) 101,762 6.74%
Mr-TyDi Zhang et al. (2021) 48,715 3.23%
DuReader He et al. (2018) 86,395 5.72%
T2Ranking Xie et al. (2023) 154,294 10.22%
Table 3: Data Statistics of E5 Dataset. We show the composition and distribution of E5 Dataset.
Dataset Task Domain #Query #Corpus Avg. D / Q
TREC-COVID Bio-Medical Bio-Medical 50 171,332 493.5
NFCorpus Information Bio-Medical 323 3,633 38.2
NQ Question Wikipedia 3,452 2,681,468 1.2
HotpotQA Answering Wikipedia 7,405 5,233,329 2.0
FiQA-2018 (QA) Finance 648 57,638 2.6
ArguAna Argument Misc. 1,406 8,674 1.0
Touché-2020 Retrieval Misc. 49 382,545 19.0
CQADupStack Duplicate-Question StackEx. 13,145 457,199 1.4
Quora Retrieval Quora 10,000 522,931 1.6
DBPedia Entity-Retrieval Wikipedia 400 4,635,922 38.2
SCIDOCS Citation-Prediction Scientific 1,000 25,657 4.9
FEVER Fact Checking Wikipedia 6,666 5,416,568 1.2
Climate-FEVER Wikipedia Fact Checking 1,535 5,416,593 3.0
SciFact Scientific Fact Checking 300 5,183 1.1
Table 4: Data Statistics of the BEIR Benchmark.
Datasets Instructions
DuReader Given a Chinese search query, retrieve web passages that answer the question.
ELI5 Provided a user question, retrieve the highest voted answers on the Reddit ELI5 forum.
FEVER Given a claim, retrieve documents that support or refute the claim.
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question.
MIRACL Given a question, retrieve Wikipedia passages that answer the question.
MrTyDi Given a question, retrieve Wikipedia passages that answer the question.
MSMARCO Passage Given a web search query, retrieve relevant passages that answer the query.
MSMARCO Document Given a web search query, retrieve relevant documents that answer the query.
NQ Given a question, retrieve Wikipedia passages that answer the question.
Squad Retrieve Wikipedia passages that answer the question.
T2Ranking Given a Chinese search query, retrieve web passages that answer the question.
TriviaQA Retrieve Wikipedia passages that answer the question.
QuoraDuplicates Given a question, retrieve questions that are semantically equivalent to the given question.
Find questions that have the same meaning as the input question.
NLI Given a premise, retrieve a hypothesis that is entailed by the premise.
Retrieve semantically similar text.
Table 5: Instructions Used for E5 Dataset.
Task Name Instruction
Arguana Given a claim, find documents that refute the claim
FiQA2018 Given a financial question, retrieve user replies that best answer the question
NFCorpus Given a question, retrieve relevant documents that best answer the question
Quora Given a question, retrieve questions that are semantically equivalent to the given question
SCIDOCS Given a scientific paper title, retrieve paper abstracts that are cited by the given paper
TREC-COVID Given a query on COVID-19, retrieve documents that answer the query
Touche2020 Given a question, retrieve detailed and persuasive arguments that answer the question
ClimateFEVER Given a claim about climate change, retrieve documents that support or refute the claim
FEVER Given a claim, retrieve documents that support or refute the claim
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question
DBPedia Given a query, retrieve relevant entity descriptions from DBPedia
NQ Given a question, retrieve Wikipedia passages that answer the question
MS MARCO Given a web search query, retrieve relevant passages that answer the query
CQADupStack Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
Table 6: Instructions Used for Evaluation on the BEIR Benchmark.
A.3 Case Study
MiniCPM This may not exactly answer your question but, as a small business owner, I would highly recommend having a professional handle your taxes. It is worth the money to have it done correctly rather than doing something wrong and getting audited or worse having penalties assessed and owing more than you thought would be possible. I would recommend this especially if this is how you make your primary income, you can always write it off as a business expense.
Debater Whether you do decide to go with a tax advisor or not, be sure to do some research on your own. When we moved to the US about 5 years ago, I did find the taxes here pretty complicated and confusing ··· After all, they are also humans prone to mistakes and your taxes are your liability in the end. My suggestion is to start with a good tool that supports tax filing for non-residents. Most of them provide a step-by-step QA based tool. As you go through the steps, Google each question you don’t understand. It may take more time than hiring a tax advisor directly but in the end it will all be worth it.
MiniCPM The novel coronavirus disease (COVID-19) pandemic is placing significant strains on health systems, scientific communities, essential public services, and economies all over the world. In this context, the worlds scientific biomedical establishment is unleashing an unprecedented response to the COVID-19 pandemic. This is a battle against time, considering the thousands of human lives are lost every day. In this commentary, based on a very recent research report, we intend to highlight how a new mechanism describing the RAGE transactivation produced by Ang II-mediated ATR1 activation can run continuously and thus, reinforcing a sustained inflammation in lungs, due to the SARS-Cov-2-mediated imbalance of the ACE/And II/ATR1 pathway.
Debater The evidence on the pathophysiology of the novel coronavirus SARS-CoV-2 infection is rapidly growing. Understanding why some patients suffering from COVID-19 are getting so sick, while others are not, has become an informal imperative for researchers and clinicians around the globe. The answer to this question would allow rationalizing the fear surrounding this pandemic. Understanding of the pathophysiology of COVID-19 relies on an understanding of interplaying mechanisms, including SARS-CoV-2 virulence, human immune response, and complex inflammatory reactions with coagulation playing a major role. An interplay with bacterial co-infections, as well as the vascular system and microcirculation affected throughout the body should also be examined. More importantly, a comprehensive understanding of pathological mechanisms of COVID-19 will increase the efficacy of therapy and decrease mortality. Herewith, presented is the current state of knowledge on COVID-19: beginning from the virus, its transmission, and mechanisms of entry into the human body, through the pathological effects on the cellular level, up to immunological reaction, systemic and organ presentation. Last but not least, currently available and possible future therapeutic and diagnostic options are briefly commented on.
Table 7: Case Studies. We present two cases from FiQA and TREC-COVID, and show the top 1 passage retrieved from MiniCPM and Debater, with the key contents in the passage highlighted.
In this subsection, we present two case studies to demonstrate the effectiveness of Debater. Specifically, we show the top 1 retrieved document comes from both MiniCPM and Debater in the FiQA and Trec-COVID datasets. The retrieval cases are shown in Table 7.
For the query “In the US, is it a good idea to hire a tax consultant for doing taxes?”, while the retrieval result of MiniCPM suggests hiring a professional, it fails to mention the specific context of “In the US”. In contrast, the document retrieved by Debater aligns more closely with the query, addressing all the relevant aspects. This highlights that Debater effectively activates the reasoning capabilities of LLMs, enabling a more fine-grained document representation.
For the query “What is the mechanism of inflammatory response and pathogenesis of COVID-19 cases?”, MiniCPM retrieves documents containing more pathological terms, such as “RAGE transactivation” and “the ACE/Ang II/ATR1 pathway”. However, these retrieved documents cannot accurately address the inflammatory response and pathogenesis of COVID-19. This suggests that MiniCPM lacks a fine-grained understanding of document content, leading to suboptimal retrieval performance. In contrast, Debater provides accurate retrieval results, demonstrating the importance of deliberate thinking before searching.
A.4 Comparison with Additional Baseline Retrievers
In this subsection, we conduct additional comparisons with baseline retrievers not discussed in the main section, including several advanced retrievers from the BEIR benchmark.
Specifically, we use CPT Neelakantan et al. (2022), Udever Zhang et al. (2023a), ULLME, Ada2, and Promptriever as our additional retrievers to compare with Debater. CPT refers to a series of large encoder models developed by OpenAI Brown et al. (2020), which are pre-trained on large-scale and unsupervised data. We use the CPT-text-175B model as our baseline model. Udever contrastively trains the BLOOM Scao et al. (2022) to build dense retrievers, enabling it to effectively generate aligned multilingual embeddings for both text and code. ULLME Man et al. (2024) enables bidirectional attention between LLMs and enhances them for text embedding using Generative Reinforcement Learning (GRL). Ada2 is a versatile text embedding model introduced by OpenAI. The evaluation result for Ada2 comes from Kamalloo et al. (2023). Promptriever Weller et al. (2025) is trained on an instance-level instruction dataset from MS MARCO, enhancing its ability to follow instructions for retrieval tasks.
The performance comparison results on the BEIR benchmark are shown in Table 8. Compared with these baseline models, Debater still exhibits competitive performance, demonstrating its effectiveness. Compared to Udever and ULLME, Debater shows significant improvement on FEVER and Climate-FEVER, indicating its effectiveness across diverse fact-checking scenarios. When compared to Promptriever, Debater demonstrates similar retrieval performance. However, unlike Promptriever, Debater does not rely on synthetic data for training its retrieval models.
Method ($\rightarrow$) CPT Udever ULLME Ada2 Promptriever Debater
Model Size ($\rightarrow$) 175B 7B 7B unk. 7B 2.4B 4B
TREC-COVID 0.649 0.838 0.836 0.813 0.839 0.795 0.836
NFCorpus 0.407 0.360 0.394 0.358 0.365 0.378 0.399
NQ / 0.533 0.614 0.482 0.619 0.560 0.561
HotpotQA 0.688 0.567 0.674 0.654 0.692 0.678 0.678
FiQA 0.512 0.367 0.423 0.411 0.459 0.434 0.462
ArguAna 0.435 0.522 0.468 0.567 0.518 0.567 0.562
Touché-2020 0.291 0.252 0.271 0.280 0.314 0.211 0.250
Quora 0.638 0.883 0.878 0.876 0.865 0.886 0.886
DBPedia 0.408 0.376 0.464 0.402 0.450 0.430 0.432
SCIDOCS / 0.189 0.211 0.186 0.173 0.197 0.212
FEVER 0.775 0.740 0.615 0.773 0.828 0.859 0.857
Climate-FEVER 0.223 0.268 0.222 0.237 0.276 0.303 0.294
SciFact 0.754 0.693 0.724 0.736 0.750 0.735 0.743
CQADupStack / 0.393 / 0.391 / 0.431 0.428
Avg CPT sub 0.525 0.533 0.543 0.555 0.571 0.571 0.582
Avg / 0.499 / / / 0.533 0.543
Table 8: Retrieval Performances of Debater and Other Baseline Models on the BEIR Benchmark.

### v2

@@ 4. Experimental Methodology
Model Size ($\rightarrow$) / 4.8B 5.8B 8B 7B / 7B 2.4B 4B
@@
Implementation Details. We initialize the Debater models with MiniCPM-2.4B and MiniCPM-4B (Hu et al., 2024). All Debater models are trained for 1,000 steps using the AdamW optimizer with a batch size of 256. The learning rate follows a cosine decay schedule, with a warm-up phase covering the first 3% of the total iterations, peaking at 2e-4. We train Debater using hybrid negatives, including one hard negative from the E5 dataset and seven in-batch negatives. The CoD length for all Debater models is set to 8. Debater is implemented using the OpenMatch toolkit (Yu et al., 2023), with flash-attention (Dao et al., 2022) and LoRA (Hu et al., 2022) enabled to mitigate memory constraints and improve computational efficiency.
5. Evaluation Results
@@
5.1. Overall Performance
The overall performance of Debater and the baseline retrievers is shown in Table 3.
5.2. Ablation Study
[…]
@@ 5.2. Ablation Study
5.3. Effectiveness of Chain-of-Deliberation with Different Thinking Depths
@@
Figure 3. Retrieval Performance of Debater with Different Thinking Depths. We set the length of the CoD to train different Debater models and evaluate them on different subsets of BEIR.
@@
Figure 4. Performance of Debater at Different Thinking Steps. We collect all documents from each thinking step to demonstrate the retrieval performance of Debater across different stages of reasoning.
5.4. Retrieval Performance of CoD-Generated Document Representations
@@
As shown in Figure 4, early steps (e.g., 1–2) produce relatively weak results, and performance may even drop.This suggests that initial embeddings, based on minimal deliberation, may lack the nuanced understanding required for effective retrieval. However, as the number of thinking steps increases, performance generally improves, indicating that more deliberation leads to more refined embeddings for retrieval. These results demonstrate that Debater leverages the CoD mechanism to refine document representations step by step by reading information from previous steps. On the other hand, the gains eventually plateau, suggesting that once embeddings become sufficiently fine-grained, further deliberation provides limited benefits while increasing computational cost.
5.5. Characteristics of the Embeddings Generated by Debater
@@
Figure 5. Similarity Relationship Between Adjacent Position Embeddings. Darker blue indicates a higher similarity score.
@@
Figure 6. Similarity Scores Between the First Seven Embeddings and the Last Embedding. The last embedding is used as the representation of documents for retrieval.
Table 5. Retrieval Performance of Debater across Different Model Scales.
Method SmolLM2-135M SmolLM2-360M MiniCPM-1B Qwen2.5-7B
Vanilla Debater Vanilla Debater Vanilla Debater Vanilla Debater
TREC-COVID 0.681 0.696 0.721 0.646 0.678 0.778 0.770 0.829
NFCorpus 0.286 0.299 0.317 0.341 0.359 0.367 0.386 0.398
FiQA-2018 0.223 0.230 0.292 0.378 0.400 0.402 0.483 0.487
ArguAna 0.455 0.428 0.480 0.481 0.548 0.534 0.561 0.577
Touché-2020 0.199 0.248 0.187 0.208 0.162 0.227 0.242 0.235
Quora 0.841 0.821 0.869 0.874 0.884 0.883 0.891 0.893
SCIDOCS 0.135 0.135 0.163 0.167 0.182 0.192 0.227 0.235
SciFact 0.530 0.554 0.576 0.674 0.704 0.705 0.763 0.767
Avg 0.419 0.426 0.451 0.471 0.490 0.511 0.540 0.553
Table 6. Case Studies. We present two cases from FiQA and TREC-COVID, and show the top 1 passage retrieved from MiniCPM and Debater. Different colors are used to annotate important content: Blue denotes critical information from the query, while Orange highlights supporting details from the passage retrieved by Debater.
Case #1 in FiQA
MiniCPM: (Less Relevant) This may not exactly answer your question but, as a small business owner, I would highly recommend having a professional handle your taxes I would recommend this especially if this is how you make your primary income, you can always write it off as a business expense.
Debater: (Most Relevant) Whether you do decide to go with a tax advisor or not, be sure to do some research on your own. When we moved to the US about 5 years ago, I did find the taxes here pretty complicated and confusing ··· After all, they are also humans prone to mistakes and your taxes are your liability in the end. My suggestion is to start with a good tool that supports tax filing for non-residents. Most of them provide a step-by-step QA based tool. As you go through the steps, Google each question you don’t understand. It may take more time than hiring a tax advisor directly but in the end it will all be worth it.
Case #2 in TREC-COVID
MiniCPM: (Less Relevant) The novel coronavirus disease (COVID-19) pandemic is placing significant strains on health systems In this context, the worlds scientific biomedical establishment is unleashing an unprecedented response to the COVID-19 pandemic. In this commentary, based on a very recent research report, we intend to highlight how a new mechanism describing the RAGE transactivation produced by Ang II-mediated ATR1 activation can run continuously and thus, reinforcing a sustained inflammation in lungs, due to the SARS-Cov-2-mediated imbalance of the ACE/And II/ATR1 pathway.
Debater: (Most Relevant) The evidence on the pathophysiology of the novel coronavirus SARS-CoV-2 infection is rapidly growing The answer to this question would allow rationalizing the fear surrounding this pandemic. Understanding of the pathophysiology of COVID-19 relies on an understanding of interplaying mechanisms, including SARS-CoV-2 virulence, human immune response, and complex inflammatory reactions with coagulation playing a major role More importantly, a comprehensive understanding of pathological mechanisms of COVID-19 will increase the efficacy of therapy and decrease mortality. Herewith, presented is the current state of knowledge on COVID-19: beginning from the virus, its transmission, and mechanisms of entry into the human body, through the pathological effects on the cellular level, up to immunological reaction, systemic and organ presentation. Last but not least, currently available and possible future therapeutic and diagnostic options are briefly commented on.
@@
Contributions of CoD Steps to Document Representations.
Figure 6 illustrates the similarity relationship between embeddings at intermediate thinking steps of CoD and the final document representation generated at the last thinking step. This helps us explore the contributions of different thinking steps to the final document representations.
@@
5.6. Effectiveness of Debater across Different Model Scales
In this subsection, we examine whether the proposed Debater remains effective across language models of varying scales.
We evaluate Debater on a range of language models with varying parameter scales, including SmolLM2-135M, SmolLM2-360M (Allal et al., 2025), MiniCPM-1B, and Qwen2.5-7B-Instruct (Yang et al., 2025). These models span from lightweight architectures suitable for edge deployment to more advanced large language models. Due to limited computational resources, our experiments were conducted on a subset of low-resource datasets from the BEIR benchmark, which still provide a diverse set of retrieval tasks for evaluation.
As shown in Table 5, the results demonstrate that Debater consistently improves average performance across diverse retrieval tasks for all model sizes. Notably, even with the smallest model, SmolLM2-135M, Debater delivers performance gains on most tasks, indicating its effectiveness under limited capacity. Moreover, as model size increases, the performance improvements become more pronounced, reflecting Debater’s ability to better leverage the increased thinking capacity and representational power of larger models. These findings highlight the robustness and scalability of Debater across LLMs of different scales.
5.7. Case Studies
In this subsection, we present two case studies on FiQA and TREC-COVID to illustrate the effectiveness of Debater. Table 6 shows the top-1 retrieved documents from MiniCPM and Debater.
For the FiQA query “In the US, is it a good idea to hire a tax consultant for doing taxes?”, MiniCPM retrieves a generic document suggesting hiring a professional but ignores the explicit “US” context. In contrast, Debater retrieves a document that fully aligns with the query, including the US-specific aspect. This shows that Debater enables finer-grained representations for contextually appropriate retrieval.
For the TREC-COVID query “What is the mechanism of inflammatory response and pathogenesis of COVID-19 cases?”, MiniCPM returns documents with many pathological terms but lacking clear explanations of the mechanisms, reflecting limited fine-grained understanding. Debater, however, retrieves results that directly address the mechanisms, demonstrating the benefit of deliberate reasoning before retrieval.
6. Conclusion
@@
Acknowledgements.
This work is partly supported by the National Natural Science Foundation of China (No. 62206042) and the Fundamental Research Funds for the Central Universities (No. N25ZLL045). This work is also supported by the AI9Stars community.